OpenStudy (anonymous):

for f(x)=x^2+8 and g(x)=x^2-6 find a) (f*g)(x) b) (g*f)(x) c. (f*g)(4)

OpenStudy (anonymous):

Is it: \[1\ f\circ g\]\[2\ f\times g\]\[3\ f\cdot g\]

OpenStudy (anonymous):

times in all of them

OpenStudy (anonymous):

Okay. Do you know how to multiply the functions?

OpenStudy (anonymous):

\[x ^{2}+8 and g(x)=x ^{2}-6\]

OpenStudy (anonymous):

yes there are three answers

OpenStudy (anonymous):

You know \(f\times g = g\times f\).

OpenStudy (anonymous):

\[(f timesx)(x)\]

OpenStudy (anonymous):

So a and b are the same here.

OpenStudy (anonymous):

b. (g times f)(x) c. (f times g)(4)

OpenStudy (anonymous):

no they are switched a is for x^2-6 b is for x^2+8

OpenStudy (anonymous):

like a. answer is x^4-12x^2+44

OpenStudy (anonymous):

i just dont know how to do the problem

OpenStudy (anonymous):

Do you know how to multiply polynomials? Like: \((x^2+8)(x^2-6)\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So that didn't work? I'm confused as to what you're struggling with.

OpenStudy (anonymous):

im having trouble of how is the answer 12x and 44

OpenStudy (anonymous):

The answer to which?

OpenStudy (anonymous):

for a x^4-12x^2+44

OpenStudy (anonymous):

Okay so what are the answers for a, b, and c and what are you getting? Did you write down the problem correctly?

OpenStudy (anonymous):

yes but i really dont know how to do them up to the end

OpenStudy (anonymous):

I'm getting for a and b, \(x^4+2x^2-48\)

OpenStudy (anonymous):

its say evaluate the exponential expression and simplify

OpenStudy (anonymous):

i came up with that to and its not right

OpenStudy (anonymous):

i\[x^4 +16x^2+58\]

OpenStudy (anonymous):

Then you must not have written the problem down correctly or something. I don't know what to tell you.

hartnn (hartnn):

if the answer is x^4-12x^2+44 then it surely is (fog)(x) and not times......

OpenStudy (anonymous):

oh thats why the times sign looks different its a o

OpenStudy (anonymous):

so how do i do that

OpenStudy (anonymous):

Remember my initial question?! Look! So it was 1 all along but you had me thinking it was 2, lol!

OpenStudy (anonymous):

im sorry

OpenStudy (anonymous):

You plug g(x) into f(x)

OpenStudy (anonymous):

yes i got tha(x^2-6)^2+8

OpenStudy (anonymous):

so \(f(g(x)) = (x^2-6)^2+8\)

OpenStudy (anonymous):

ok so do we times -6 with the power of 2 to get 12

OpenStudy (anonymous):

\(f(g(x)) = (x^2-6)^2+8= (x^2-6)(x^2-6)+8\)

OpenStudy (anonymous):

We have to use foil here to expand the \((x^2-6)^2\)

OpenStudy (anonymous):

x^4+12x+36

OpenStudy (anonymous):

then add 8

OpenStudy (anonymous):

?

OpenStudy (anonymous):

It should be \(-6x-6x = -12x\) not \(+12x\)

OpenStudy (anonymous):

oh duh -12x^2

OpenStudy (anonymous):

its midnight here and i havent slept in so long

OpenStudy (anonymous):

ok i have the answer to b. x^4+16x^2+58

OpenStudy (anonymous):

c. is 138

OpenStudy (anonymous):

omgosh thank you